Company x plans to increase the work force by 20% to a total of 8460 employees. How many people does company X employ?
Answer by macston(5194)

.
Another way to phrase the question: 8460 is 120% of what number?
.
1860/1.20=7050
ANSWER: The company employs 7050 people.
.
CHECK:
.
20% 0f 7050=(0.20)(7050)=1410
7050+1410=8460, the total of the planned increase.